A man travels 7.0km due North,then 10.0km East,find the resultant displacement.If he travels 7.0km 30degrees East of North,then 10.0km East,find the resultant displacement
1
Expert's answer
2019-10-28T11:36:31-0400

1)


"d=\sqrt{7^2+10^2}=12.2\ km"

"\theta=\arctan{\frac{7}{10}}=35\degree \ North\ of \ East"

2)


"d=\sqrt{7^2+10^2+2(7)(10)\cos{(90-30)}}=14.8\ km"

"\theta=\arctan{\frac{7\cos{30}}{10+7\sin{30}}}=24.2\degree \ North\ of \ East"